OpenStudy (anonymous):

The numbers 7, 24, and c are a Pythagorean triple. What is the value of c?

OpenStudy (anonymous):

25=c

OpenStudy (anonymous):

7^2+24^2=c^2 625=c^2 25=c

OpenStudy (anonymous):

thanks @heath8280